Juliana is considering an investment proposal with the following cash flows:

Initial investment-depreciable assets $36,000
Net cash inflows from operations (per year for 10 years) 6,000
Disinvestment 0

For parts b. and c., round answers to three decimal places, if applicable.

a. Determine the payback period.

b. Determine the accounting rate of return on initial investment.

c. Determine the accounting rate of return on average investment.
Answer #1

Solution a:

Payback period = Initial investment / Annual cash inflows = $36,000 / 6000 = 6 years

Solution b:

Annual net income = Annual cash inflows - Depreciation = $6,000 - ($36,000 / 10) = $2,400

Accounting rate of return = Average annual income / Initial investment = $2,400 / $36,000 = 6.667%

Solution c:

Average investment = (Cost + Salvage value) / 2 = ($36,000 + 0) / 2 = $18,000

Accounting rate of return = Average annual income / Average investment = $2,400 / $18,000 = 13.333%